OpenStudy (anonymous):

A screw jack has 90 threads to the metre; the effort is applied at the end of an arm 30 cm long. What force must be applied to lift a load of 120 kg, if the efficiency at that load is 25% ?

OpenStudy (whpalmer4):

F = (Q · p) / (2 · π · R) F is the force applied Q is the load p is thread pitch (distance for 1 turn) R is lever arm radius Here we have \(Q = 120 kg\), \(p = 1/90 m\), \(R = 0.3 m\) and our net force will be quadrupled because the efficiency is only 25%.